Professional Products Inc., a wholesaler of office products, was organized on February 5 of the current year, with an authorization of 26,000 shares of preferred 2% stock, $100 par and 550,000 shares of $10 par common stock. The following selected transactions were completed during the first year of operations:

Feb. 5 Issued 247,000 shares of common stock at par for cash.
5 Issued 300 shares of common stock at par to an attorney in payment of legal fees for organizing the corporation.
Apr. 9 Issued 69,100 shares of common stock in exchange for land, buildings, and equipment with fair market prices of $153,500, $488,060 and $160,000 respectively.
June 14

Issued 15,000 shares of preferred stock at $109 for cash .MAKE A JOURNAL
Answer #1

Journal

February 5 Cash 2,470,000
Common stock 2,470,000
February 5 Organization expense 3,000
Common stock 3,000
April 9 Land 132,328
Building 420,741
Equipment 137,931
Common stock 691,000
June 14 Cash 1,635,000
Preferred stock 1,500,000
Paid in capital, in excess of par - Preferred 135,000
Fair market value Percentage of Fair market value Value to be recorded
Land 153,500 153,500/801,560 = 19.1501572% 691,000 x 19.1501572% = 132,328
Building 488,060 488,060/801,560 = 60.8887669% 691,000 x 60.8887669% = 420,741
Equipment 160,000 160,000/801,560 = 19.9610759% 691,000 x 19.9610759% = 137,931
801,560 691,000

Please ask if you have any query related to the question. Thank you